OpenStudy (anonymous):

use the second equation to make a substitution for y in the first equation. 3x + y = 1 y + 4 = 5x What is the resulting equation?

OpenStudy (anonymous):

x=5/8

OpenStudy (anonymous):

thank you katie

OpenStudy (anonymous):

anytime :)

OpenStudy (jessicajohnson):

y= (5x-4) 3x+(5x-4)=1 8x-4=1 8x=5 x=5/8

OpenStudy (anonymous):

thank you jessica :D